HUBBLE SPACE TELESCOPE DAILY REPORT #1674 PERIOD COVERED: 0000Z (UTC) 07/12/96 - 0000Z (UTC) 07/15/96 Daily Status Report as of 197/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Five Sets of WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with flu= x excesses at 3000 A0 and galactic sources with excesses at 1600 A0. The = observations were completed as planned, and no problems were reported. 1.2 Completed GHRS/1/FOS/BL 6678 (Nearby Young Dwarf Galaxies and = Their Lyman Alpha Emissions) The High Resolution Spectrograph (Side One) and the Faint object = Spectrograph (Blue Side) were used to make observations of the young, = metal deficient, blue, compact, dwarf galaxy TOL65 in order to study its = Lyman Alpha emissions. The observations were completed as planned, and = no problems were reported. 1.3 Completed FOS/RD 6925 (FOS Observations of Solar Analog Standards) The Faint Object Spectrograph (Red Side) was used to make absolute = spectrophotmetry observations of the solar analog star P177D-CAlI in the = spectral range of 2200 A0 to 8500 A0. The objective is to provide the = primary standards for the absolute calibration of the NICMOS. The = observations were completed as scheduled, and no problems were reported. 1.4 Completed FOC/96 6197 (FOC Cycle 5 Polarization PSF = Characterization) The Faint Object Camera (f/96) was used to make observations of the = globular cluster Omega Cen in order to determine the FOCs point spread = function polarization characterization. The FOCs point spread function = will be determined from a series of dithered images made of a = non-polarized point source within the globular cluster. Numerous = different filters were used. During this proposal, the early load of the = DF-224 load occurred (described in Section 2.3). The full effect of this = anomalous event are not currently known. 1.5 Completed Three Sets of WF/PC-2 5971 (HST Medium Deep Survey) The WF/PC-2 was used to make a long exposure medium deep sky = survey observations at latitudes larger than 10 degrees taken in parallel = with the prime science observations. During the second iteration of this = proposal, the early load of the DF-224 load occurred (described in = Section 2.3). The full effect of this anomalous event are not currently = known. The other iterations were executed as scheduled, and no anomalies = were noted. 1.6 Completed WF/PC-2 6361 (High Resolution Imaging of X-Ray Selected = AGNs) The WF/PC-2 was used to make high resolution observations of the = known low red shifted x-ray active galactic nuclei (AGN) MS0412.4-0 and = MS0731.6+8. These data will allow a study of the optical properties of = an unbiased sample of AGN host galaxies. The observations were completed = as scheduled, and no problems were reported. 1.7 Completed Two Sets of FOS/RD 6577 (An Expanded Survey for Study of = the Lyman-Alpha Line and QSO Mg II Absorption) The Faint Object Spectrograph (Red Side) was used to make = observations of the galaxies PHL938 and PKS0402-13 in order to make a = survey of the Lyman-Alpha line and the 68 Mg II absorption line system. = The observations were completed as scheduled, and no problems were = reported. 1.8 Completed WF/PC-2 6348 (Blue Continuum Snapshots of 3CR Radio = Galaxies) The WF/PC-2 was used to make some snapshot images of the blue = continuum in radio galaxy 3C288.1. Combined with existing data, these = exposures will provide high spatial resolution information on the color = distribution within galaxies, redding due to dust, spacial indices of the = optical synchrotron jets, and colors of the active nuclei and tidal jets. = The observations were executed as scheduled, and no anomalies were = noted. 1.9 Completed FOS/RD 6585 (UV Spectra of Representative Stellar = Populations of Age > 2) The Faint Object Spectrograph (Red Side) was used to make = mid-ultraviolet spectrographic observations of nuclear region of nearby = galaxies NGC-6127 and NGC-7619. These galaxies have a relatively old = stellar population. The objective is to fill in crucial gaps in our = understanding of stellar populations, and thus aid in the modeling of = galactic evolution. The observation was completed as planned, and no = anomalies were reported. However, during the second iteration, the = safemode entry described in Section 2.3 occurred about five hours into = this nine-hour proposal. 1.10 Completed WF/PC-2 6250 (WF/PC-2 Cycle 5 Calibration: Internal = Monitor) The WF/PC-2 was used to make a series on internal calibration = observations. Internal flats, biases, and K-Spot calibration images were = taken. These observations are designed to monitor the health of the = cameras. The observations were executed as scheduled, and no anomalies = were noted. The following scheduled proposals were not executed due to the safemode = entry described in Section 2.3: WF/PC-2 6253 (Pure Parallel Archive Program) WF/PC-2 6361 (High Resolution Imaging of X-Ray Selected AGNs) Two Sets of WF/PC-2 6187 (WF/PC-2 Cycle 5 Calibration: Earth Flat) FOC/96 6349 (Snapshot Search for High Redshift QSOs with Far-UV Flux) GHRS/1/2 6410 (Low Red Shift Lyman Alpha Clouds) GHRS/1/2 6594 (White Dwarf Companions of Barium and CH Peculiar Stars) WF/PC-2 6749 (A Study of the Gaseous Halo of the Andromeda Galaxy) GHRS/1/2 6413 (High Velocity Thermal Shock Instabilities in the Vela Supernova Remnant) 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 17 Successful : 17 2.2 FHST Updates: Scheduled : 21 Successful : 21 2.3 Operations Notes: At 194/160111Z, the DF-224 load was uplinked early. Operations = personnel were able to restore normal operations, but some portions of = FOC 5971 and WF/PC-2 6197 were affected. The spacecraft entered Inertial Hold safemode at approximately = 195/1731Z as a result of failing the Earth/Moon Bright Object test. This = failure was the result of a step change (bit flip of the 18th bit of the = higher order word) in the second element of the vehicle attitude = quaternion that occurred at 195/164635Z. Parity did not fail, although = the bit flip had the appearance of an SEU. No other anomalous event have = been identified. HSTAR 5656 was written. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH